Corso di Matematica - Scienze geologiche - 2003-04 Risultati del primo compitino

matricola nome  voto matricola nome  voto
ANONIMO  9  Limoncelli  17
Adami  11  Loiaconi  7
Altomonte  15  Maggiore  15
Anfuso  24  Mancuso  7
Anzilotti  22  Manzi  13
Babboni  11  Marinari  29
Banci  10  Merciadri  10
Beconcini  18  Nencini  11
Bendinelli  21  Nicolai  10
Bertoncini  19  Nicolardi  14
Biagioni  15  Orsetti  27
Bonaventuri  7  Orlandini  30
Borzoni  14  Pitanti  10
Brandani  9  Puccetti  19
Calò  13  Puccini  6
Campinoti  9  Renucci  13
Camposano  19  Ridolfi  25
Ciuffi  9  Riva  2
Curcio  26  Rosa  22
Dagani  21  Rossi  21
Damiani  11  Salcioli  19
De Filippo  25  Savio  15
Duranti  7  Scattareggia  6
Garzelli  17  Trunfio  17
Grossi  3  Vagli  15
Iuliano  18  Ventre  22
Landi  20
Libardi  11